Grade 9General Physics

a object cover 20 cm in 1 second and in 5 second object cover 180 cm .find acceleration

Dhairya Jain
9 Years agoGrade 9

1 Answer

Arnab Nanda Goswami
9 Years ago
From the equation of linear motion – [ S=ut+1/2at] we get.......
                  20=ux1 + 0.5ax1(from 1st information)
            Or,     2u + a=40.......(i)
Now.........
                     180=ux5 + 0.5ax5(from 2nd information)
             Or,    2u + 5a=72.......(ii)
By solving both the equations ue get.........
                      4a=32
             Or,     a=8 m/t2
Ans: the acceleration  of the object is 8 m/t2